OpenStudy (sarahc):

Find a ⋅ b. a = <2, 4>, b = <2, 5> <4, 9> 24 <4, 20> 16

OpenStudy (jdoe0001):

\(\bf <a,b>\cdot <c,d>\implies (a\cdot c)+(b\cdot d)\impliedby \textit{dot product}\)